For f(x) = x - 3 and g(x) = x2 +1,

Need help finding " (g/f)(x) " and " (g ° f)(x) "
Thank youcfor your time.
Answer #1

f(x) = x-3 and g(x) = x²+1 ( )x) and (gof)(x) ☺ (9)(x) = g(x) f(x) x 71 X-3 (31 x² X-3 x) @ (90+)(x) = g(f(x)) 9 (x-3) (x - 3